OpenStudy (anonymous):

You are given the identity x^2 -12x+10 ≡(x+a)^2 + b Work out the values of a and b.

OpenStudy (anonymous):

Try expanding the right side using foil (x+a)^2=(x+a)(x+a)

OpenStudy (anonymous):

x^2-12x+36-36+10 (x-6)^2-26 a=6 b=-26

OpenStudy (anonymous):

Nevermind, that's a much better way of doing it! @Kimthoanguyen used completing the square, very clever.

OpenStudy (anonymous):

I missed it: a=-6 b=-26